# SymCode: A Neurosymbolic Approach to Mathematical Reasoning via Verifiable Code Generation

### v1

@@ 6. Print ONLY the final answer in LaTeX boxed form:
How many distinct values can be obtained from the expression
$2\cdot 3\cdot 4\cdot 5+1$ by inserting any number of parentheses?
@@
In $\triangle ABC$, $AB=4,BC=6,AC=8$. Squares $ABQR$ and $BCST$ are
drawn external to the triangle. Compute the length of QTQT.
@@
Let $\triangle ABC$ have circumcenter $O$ and incenter II with $\overline{IA}\perp\overline{OI}$,
circumradius 1313, and inradius 66. Find $AB\cdot AC$.
@@ 4.1 Datasets
We evaluate SymCode on three widely used, challenging benchmarks that require multi-step mathematical reasoning, spanning difficulty from high school competitions to Olympiad-level problems: (1) MATH-500, a 500-problem subset of the MATH dataset, comprising challenging problems from high school mathematics competitions, covering topics like algebra, geometry, number theory, and precalculus Lightman et al. (2023); (2) OlympiadBench, 674 text-only English math problems from national and international olympiads requiring creative and formal reasoning He et al. (2024); and (3) American Invitational Mathematics Examination (AIME) 2024 & 2025, 60 (30 from each year) short-answer problems from recent competitions that bridge high school and olympiad difficulty Mathematical Association of America (2024). Table˜2 shows a sample from each dataset.
@@ 5.3 Token Efficiency Analysis
Consistent with our initial hypothesis, SymCode is substantially more token-efficient than its prose-based counterparts. Python code expresses complex operations concisely, whereas natural language requires verbose descriptions. On average, SymCode solutions used just 699 output tokens, proving significantly more concise than the baselines: Tree of Thoughts (1770 tokens), Decomposition (1962 tokens), and Chain of Thought (2991 tokens). This results in a token reduction of approximately 60% to 77% compared to these prose-based methods, a significant advantage for inference cost and latency. The self-debugging loop in SymCode+ raises the average token count to 890. This overhead is most pronounced for Llama 3.2, whose higher self-debugging activation rate requires more token-intensive refinement.
@@ 5.4 Qualitative Error Analysis
In contrast, SymCode shifts failure modes toward the structured stage of problem setup, with most errors due to problem misinterpretation (56.2%) and incorrect API usage (31.3%). The final 12.5% of “Other” errors are mostly runtime issues like infinite loops or verification failures where an assertion check fails. This transition from opaque reasoning errors to transparent, programmatic ones points to clearer paths for improvement through better code generation and debugging. For a direct, side-by-side comparison illustrating how a baseline method and SymCode fail on the same problem, see Appendix B.
@@ 29 C_sol = solutions[0]
Appendix B Qualitative Failure Analysis Examples

### v2

@@ 6. Print ONLY the final answer in LaTeX boxed form:
How many distinct values can be obtained
from the expression $2\cdot 3\cdot 4\cdot 5+1$ by
inserting any number of parentheses?
@@
In $\triangle ABC$, $AB=4,BC=6,AC=8$.
Squares $ABQR$ and $BCST$ are drawn external
to the triangle. Compute the length of QTQT.
@@
Let $\triangle ABC$ have circumcenter $O$ and incenter
II with $\overline{IA}\perp\overline{OI}$, circumradius 1313, and
inradius 66. Find $AB\cdot AC$.
@@ 4.1 Datasets
We evaluate SymCode on three widely used, challenging benchmarks that require multi-step mathematical reasoning, spanning difficulty from high school competitions to Olympiad-level problems: (1) MATH-500, a 500-problem subset of the MATH dataset, comprising challenging problems from high school mathematics competitions, covering topics like algebra, geometry, number theory, and precalculus Lightman et al. (2023);
(2) OlympiadBench, 674 text-only English math problems from national and international olympiads requiring creative and formal reasoning He et al. (2024); and
(3) American Invitational Mathematics Examination (AIME) 2024 & 2025, 60 (30 from each year) short-answer problems from recent competitions that bridge high school and olympiad difficulty Mathematical Association of America (2024). Table˜2 shows a sample from each dataset.
@@ 5.3 Token Efficiency Analysis
Consistent with our initial hypothesis, SymCode is substantially more token-efficient than its prose-based counterparts. Python code expresses complex operations concisely, whereas natural language requires verbose descriptions. On average (across all three datasets and all three models), SymCode solutions used just 699 output tokens [^1] See Appendix B for the token-counting protocol and model-specific token visibility. All reported token counts are rounded., proving significantly more concise than the baselines: Tree of Thoughts (1770 tokens), Decomposition (1962 tokens), and Chain of Thought (2991 tokens). This results in a token reduction of approximately 60% to 77% compared to these prose-based methods, a significant advantage for inference cost and latency. The self-debugging loop in SymCode+ raises the average token count to 890. This overhead is most pronounced for Llama 3.2, whose higher self-debugging activation rate requires more token-intensive refinement.
@@ 5.4 Qualitative Error Analysis
In contrast, SymCode shifts failure modes toward the structured stage of problem setup, with most errors due to problem misinterpretation (56.2%) and incorrect API usage (31.3%). The final 12.5% of “Other” errors are mostly runtime issues like infinite loops or verification failures where an assertion check fails. This transition from opaque reasoning errors to transparent, programmatic ones points to clearer paths for improvement through better code generation and debugging. For a direct, side-by-side comparison illustrating how a baseline method and SymCode fail on the same problem, see Appendix C.
@@ 29 C_sol = solutions[0]
Appendix B Token Accounting Details
This appendix describes how the token-efficiency numbers in Section 5.3 are computed.
We count output tokens produced by the LLM (i.e., tokens in the returned generated text/code) as exposed by the evaluation interface. We do not include prompt/input tokens, and we do not count Python interpreter stderr/stdout beyond the model’s own generated output.
Model-specific visibility of internal reasoning tokens.
  • GPT-5-nano: Only returned response/completion tokens were available in our setup; internal/hidden reasoning tokens were not exposed. Thus, for GPT-5-nano, the reported token counts reflect visible output tokens for all methods (SymCode and prose baselines).
  • Llama 3.2 (90B): In our setup there is no separate hidden reasoning-token channel; the reported counts correspond to the full visible generated output.
  • GPT-OSS (20B): Internal reasoning tokens are observable in the open-source setting; we include them in the token count for all methods under GPT-OSS.
Each evaluated run (one problem instance $\times$ one model $\times$ one prompting method) yields one output-token count. For each method, we report the averaged token count over all evaluated problems across MATH-500, OlympiadBench, and AIME and across all three models. All values are rounded to the nearest integer for readability.
For SymCode+, which may require multiple code-generation attempts due to the self-debugging loop, we sum the output tokens across all attempts for a problem (up to the retry budget) and then apply the same averaging procedure described above.
Appendix C Qualitative Failure Analysis Examples